accounting

Cullumber Company makes 2 products, sleep masks and sleep socks. Additional information follows:



Sleep Masks


Sleep Socks


Units


6,500

5,000

Sales


$135,000

$100,000

Variable costs


61,030

32,000

Fixed costs


34,320

34,000

Net income


$39,650

$34,000

Profit per unit


$6.10

$6.80


Compute the Contribution margin per unit. (Round answers to 2 decimal places, e.g. 15.25.)



Contribution margin per unit


Sleep Masks


$enter a dollar amount rounded to 2 decimal places 

Sleep Socks


$enter a dollar amount rounded to 2 decimal places 


Cullumber has unlimited demand for both products. Therefore, which product should the company encourage its sales people to emphasize?

select a product
Answer #1

The formula to compute for contribution margin per unit is:


Contribution Margin Per Unit = (Sales - Variable Cost) / Units Sold

Sleep Mask = (135,000 -61,030) / 6,500 = 11.38

Sleep Sock = (100,000 -32,000) / 5,000 = 13.60



Contribution Margin per Unit
Sleep Masks                                  11.38
Sleep Socks                                  13.60


Based on the above table, Cullumber Company should focus on selling sleep socks as it has higher contribution margin than sleep masks.

answered by: Eunise Tumambing